# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v3

@@ Abstract
We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities. Inspired by covariant theory in physics, we address this problem by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this foundation, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of elements and challenging scenarios show dramatic breakthroughs on the state-of-the-art prediction accuracy, with improvements of up to 40% in predicting Hamiltonians and up to 76% in predicting downstream physical quantities such as occupied orbital energy. Our approach goes beyond handling physical systems and offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm.
keywords:
Deep Learning, Neural Networks, SO(3)-equivariance, Atomic Modeling, Hamiltonian Prediction
@@ 1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research [1]. For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7, 8], accelerating processes by thousands or even millions of times. This reduction in computational demand not only unlocks possibilities for analyzing extremely large atomic systems but also paves the way for efficient material simulation, discovery, and design that were previously unimaginable.
@@
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Traditional deep learning methods [9, 10] typically approximated SO(3)-equivariance through rotational data augmentation. However, this approach does not provide the strict assurance of SO(3)-equivariance required for physical research. Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, existing deep learning methods are unable to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. As pointed out in previous literature [11], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance. Although several efforts have been made to alleviate this issue by introducing gated activation mechanisms [8, 7, 12, 13, 14], developing local coordinate strategy [6], and employing discrete spherical channels [11, 15, 16], they still, to varying degrees, compromised either non-linear expressiveness [7, 8, 12, 13, 14] or strict SO(3)-equivariance [11, 15, 16] when processing physical systems. Please refer to Supplementary Information A for an in-depth analysis.
@@
To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity [18], the spacetime interval between events is an invariant under Lorentz transformations. This invariance is foundational in formulating physical laws that involve covariant quantities. Furthermore, from the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities. Built upon these insights, we theoretically solve the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations from these SO(3)-invariant ones. Specifically:
@@
First, we construct quantities that are theoretically SO(3)-invariant, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $\mathbf{Q}$ are the SO(3)-equivariant regression targets, $tr(\cdot)$ signifies the trace operation, and $\dagger$ denotes the conjugate transpose operation. A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
@@
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$, from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
@@
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as shown in Fig. 1. Our method dramatically enhances the state-of-the-art performance on eight benchmark Hamiltonian databases from the renowned DeepH and QH9 benchmark series [6, 8, 19], showing excellent generality to various challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories, with an accuracy improvement of Hamiltonians of up to 40%. Moreover, our approach substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonians including occupied orbital energies and electronic wavefunctions, with an accuracy improvement by a maximum of 76% and 17%, respectively. Our leading performance comprehensively demonstrate that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
Figure 1: Illustration of the high-dimensional complexity of the Hamiltonian. In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs. Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 2 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here. Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
@@ 2 Problem Formalization
Physical quantities that are SO(3)-equivariant typically possess their respective equivariant formulas under both matrix-formed direct-product states and vector-formed direct-sum states, which are mathematically equivalent to each other.
Without loss of generality, let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$. It obeys the following SO(3)-equivariant law:
[…]
where $\dagger$ denotes the conjugate transpose operation. ${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix and ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$, ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively, ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components with vector forms, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq. (1) but with a simpler form:
@@
where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respective denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
Eq. (1) and Eq. (2) are mathematical equivalent to each other, meaning that fulfilling one under the corresponding state naturally satisfies the other. Due to the simplicity and ease of handling of the direct-sum state form, existing deep learning methods [6, 8, 7] typically first predict $\mathbf{q}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, then convert it to $\mathbf{Q}^{l_{p}\otimes l_{q}}$ through Clebsch-Gordan matrices. In the direct-sum state, to obey the equivariant law of Eq. (2) for the regression target, the hidden representations of neural networks must also satisfy the same form of equivariance:
@@
where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings denoted as $g_{nonlin}(\cdot)$ to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance. Meanwhile, $g_{nonlin}(\cdot)$ must also preserve SO(3)-equivariance, which is expressed as:
${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $Silu$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, is the core problem this paper aims at solving.

### v4

@@ Abstract
We propose both a theoretical and a methodological framework to address a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities. Inspired by covariant theory in physics, we present a solution by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove, establishing a strong foundation for regressing complex SO(3)-equivariant targets. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of systems and challenging scenarios show substantial improvements on the state-of-the-art prediction accuracy of deep learning paradigm. Our method boosts Hamiltonian prediction accuracy by up to 40% and enhances downstream physical quantities, such as occupied orbital energy, by a maximum of 76%. Our method also significantly promotes the acceleration ratios for the convergence of traditional Density Functional Theory (DFT) methods.
@@ 1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research (Zhang et al., 2023). For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ (Unke et al., 2021; Gu et al., 2022; Li et al., 2022; Yu et al., 2023b; Gong et al., 2023), opening up new possibilities for analyzing extremely large atomic systems, enabling efficient materials simulation and design, as well as groundbreaking molecular pharmaceutical research that was previously unimaginable.
@@
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, it is challenging for deep learning methods to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy on modeling physical systems. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. This issue is commonly found in physics-oriented machine learning tasks that demand both strict equivariance and fine-grained generalization performance, as analyzed by Zitnick et al. (2022). Several recent efforts have been made to alleviate this issue (Zitnick et al., 2022; Passaro & Zitnick, 2023; Wang et al., 2024c; Yin et al., 2024), however, they often compromised strict SO(3)-equivariance.
@@
To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity (Resnick, 1991), the spacetime interval between events is invariant under Lorentz transformations, and is fundamental in formulating physical laws that involve equivariant quantities. For another example, the total electronic energy of an atomic system is SO(3)-invariant, while the derived force fields are SO(3)-equivariant. We aim to extend the relationship between invariance and equivariance from specific physical quantities to a more general representation learning framework. From the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities. Built upon these insights, we propose a solution to the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations and the target quantities from these SO(3)-invariant ones. Specifically:
@@
First, we propose a theoretical construct of SO(3)-invariant quantities, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $tr(\cdot)$ signifies the trace operation, $\dagger$ denotes the conjugate transpose operation, and $\mathbf{Q}$ denotes the SO(3)-equivariant regression targets we aim to predict. A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
@@
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$ in the inference phase from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ in the training phase. Taking SO(3)-invariant features as a bridge, this mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove, laying a solid foundation for accurately inferring complex SO(3)-equivariant targets.
@@
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deep understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as pointed out by Yin et al. (2024). Our method significantly improves the state-of-the-art performance in Hamiltonian prediction on eight databases from the well-known DeepH and QH9 benchmark series (Li et al., 2022; Gong et al., 2023; Yu et al., 2023a). It demonstrates excellent generalization performance to both crystalline and molecular systems, covering challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories. Furthermore, as observed from the experiments on the QH9 benchmark series, our approach also substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonian including occupied orbital energy and electronic wavefunction. Moreover, our method also demonstrates superior performance in accelerating the convergence of classical DFT by providing predicted Hamiltonians as initialization matrices. Our leading performance comprehensively demonstrates that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
2 Related Work
The SO(3)-equivariant representation learning paradigm (Thomas et al., 2018; Geiger & Smidt, 2022) typically developed group theory-based symmetry operators, such as linear scaling, element-wise sum, direct products, direct sums, Clebsch-Gordan decomposition, and equivariant normalization, to encode equivariant features. These operators have been used to construct graph neural network architectures for tasks in 3D point cloud analysis (Fuchs et al., 2020), molecular property prediction and dynamic simulation (Musaelian et al., 2023; Liao & Smidt, 2023), as well as Hamiltonian prediction (Schütt et al., 2019; Unke et al., 2021; Gong et al., 2023; Zhong et al., 2023). However, as non-linear activation functions may result in the loss of equivariance, they are restricted when applied to SO(3)-equivariant features with degree $l$ greater than one. This restriction severely limits the network’s capability to model complex non-linear mappings. To alleviate this issue, methods like DeepH-E3 (Gong et al., 2023), QHNet (Yu et al., 2023b) and Equiformer (Liao & Smidt, 2023) introduced a gated activation mechanism that feeds SO(3)-invariant features ($l=0$) into non-linear activation functions and uses these features as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. Nevertheless, multiplying SO(3)-equivariant features with linear coefficients may not fundamentally improve their non-linear expressiveness.
In order to improve non-linear expressiveness, Zitnick et al. (2022) decomposed SO(3)-equivariant features into SO(3)-invariant coefficients of spherical harmonic basis functions. These SO(3)-invariant coefficients were processed by non-linear neural networks to enhance expressiveness, with equivariance regained by recombining the updated coefficients with the basis functions. In subsequent developments (Passaro & Zitnick, 2023; Liao et al., 2024; Wang et al., 2024b; c), this approach has demonstrated a remarkable capacity to fit complex functions. However, as pointed out by existing literature (Zhang et al., 2023), this approach degenerates from continuous to discrete rotational equivariance, losing strict equivariance to continuous rotational transformations due to the decomposition based on inner-product operations with discrete basis functions; Li et al. (2022) proposed a local coordinate strategy, projecting rotating global coordinates onto SO(3)-invariant local ones for the non-linear neural network to encode. However, this strategy is effective only for global rigid rotations and fails to maintain symmetry under non-rigid perturbations like thermal fluctuations or bilayer twists, as it lacks a neural mechanism to enforce equivariance; Yin et al. (2024) proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians. In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy. However, the equivariance achieved through data-driven methods does not have strict theoretical guarantee, even with rotational data augmentation. In some applications, the demands for symmetry are extremely high. Even minor deviations from perfect SO(3)-equivariance can result in incorrect physical results. In this paper, we propose a framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks to resolve the equivariance-expressiveness dilemma.
@@ 3 Problem Formalization
For a background introduction on groups, group representations, equivariance, and invariance, please refer to Appendix A. Here, we directly focus on the equivariance of physical quantities under 3D rotational operations that form the SO(3) group. Let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$. It obeys the following SO(3)-equivariant law:
${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}={\bf{D}}^{l_{p}}({\bf{R}})\cdot{\bf{Q}}^{l_{p}\otimes l_{q}}\cdot{({\bf{D}}^{l_{q}}({\bf{R}}))^{\dagger}}$ (1)
where $\dagger$ denotes the conjugate transpose operation. ${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix, ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$ and ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively; ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq. 1 but with a simpler form:
@@
where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respectively denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
For ease of processing, the internal representations of SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023) are typically in the direct-sum form. To obey the equivariant law of Eq. 2 for the regression target, these hidden representations must also satisfy the same form of equivariance:
@@
where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of hidden features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance. Meanwhile, the non-linear mappings, denoted as $g_{nonlin}(\cdot)$, must also preserve SO(3)-equivariance, which is expressed as:
${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},\,\,\,\,\text{subject to}\,\,{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $SiLU$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, and effectively apply it to the prediction of SO(3)-equivariant complex physical quantities, i.e., electronic-structure Hamiltonian in this context, is the core problem this paper aims at solving.